# for preordered semifields and semirings II

### v4

@@ 5.  Type classification of multiplicatively Archimedean
By reversing the order and replacing $x$ by $x^{-1}$ and $r$ by $1-r$, the second inequality reduces to the first. We therefore only prove the first.
Since the expression $rx+(1-r)$ is obviously non-strictly monotone in $r$, it is enough to prove the claim for $r=2^{-n}$ with $n\in\mathbb{N}_{>0}$, in which case it amounts to
@@
    which is equivalent to the desired inequality with $\ell^{\prime}$ in place of $\ell$. Therefore it is enough to show that if the statement holds for given $\ell\in\mathbb{N}$, then it also holds for $12\ell+$ 1.
    This step from $\ell$ to $12\ell+$ 1 works as follows,
@@
Now if $\leq$ was arctic, then we would have $x_{i}+y_{i}\sim x_{i}y_{i}+1$ by Lemma 5.12, and therefore $a+z\approx 1+z$ with $z\coloneqq\sum_{i=1}^{\ell}c_{i}(x_{i}y_{i}+1)+d$. Upon adding a quasi-complement of $z$ on both sides, we obtain further $a+n\approx 1+n$ for some $n\in\mathbb{N}$. By chaining, we can reduce to the case $n=1$ without loss of generality. But then applying the cancellation criterion of Proposition 5.10 to $11+a\leq 1+$ 1 shows $a\leq 1$, contradicting the initial assumption $a>1$. ∎
@@ 6.  The ambient preorder
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    What we need to prove is that $44u+u^{-1}>>$ 4, which unfolds to
@@
[…]
@@
As the proof will show, the assumption on $(F/\!\sim)\otimes\mathbb{Z}$ can in fact be weakened to formal smoothness over $\mathbb{Q}$ (and probably further). We nevertheless phrase the statement in terms of the stronger assumption stating that this ring should be a finite product of fields, since this is how we will use the statement later and is more elementary.
@@
[…]
@@
[…]
@@
is an ideal in $R\coloneqq(F/\!\approx)\otimes\mathbb{Z}$, namely precisely the kernel of the canonical projection homomorphism $(F/\!\approx)\otimes\mathbb{Z}\longrightarrow(F/\!\sim)\otimes\mathbb{Z}$. We have $J^{2}=0$, since $a\sim b$ and $c\sim d$ imply $(b-a)(d-c)=0$ in $R$ via Lemma 5.12 and
@@
assuming $a,c\neq 0$ without loss of generality. Hence $R$ is a square-zero extension of the quotient ring $R/J\cong\prod_{i=1}^{n}K_{i}$. Each $K_{i}$ is formally smooth over $\mathbb{Q}$ [8, Corollary 9.3.7], and therefore also their product is formally smooth over $\mathbb{Q}$. (A finite product of formally smooth algebras is formally smooth by lifting of idempotents.) In particular, the square-zero extension $R\to R/J$ is split by a ring homomorphism $\beta:R\to R/J$. In the semialgebra case, $R$ is an $\mathbb{R}$-algebra in a canonical way, and in this case we have formal smoothness over $\mathbb{R}$ for the same reason, so that we can choose $\beta$ to be $\mathbb{R}$-linear.
The universal property of $(F/\approx)\otimes\mathbb{Z}$ implies that $\phi:F\to\mathbb{R}_{+}$ uniquely extends to a ring homomorphism $\hat{\phi}:R\to\mathbb{R}$. In the following, we will construct a $\mathbb{Q}$-linear map
@@
    Indeed, writing $a=c-b$ for $b\sim c$ in $F^{\times}$ and plugging in the definition of $\hat{D}$ shows that this holds for all $r\in F^{\times}$. But this is nough by linearity in $r$ since $R=F^{\times}-F^{\times}$.
@@
It is interesting to ask how unique the derivation $D$ constructed in the proof is. Clearly $D$ can be replaced by any positive multiple ot itself, but is there more freedom? We answer this question now.
@@ 7.  A stronger catalytic Vergleichsstellensatz
We start the proof with the case where $S$ is a preordered semifield $F$ of polynomial growth. In this case, the two additional hypotheses amount to $F$ having quasi-complements and that $F/\!\sim\mathop{\otimes}\mathbb{Z}$ is a finite product of fields. We fix a power universal element $u\geq 1$ in $F$ and also assume $y=1$ without loss of generality. Considering this case will take the bulk of the proof; we will generalize from there in the final two paragraphs.
If $x\not\leq 1$, then by Theorem 2.1 we can find a total semifield preorder $\preceq$ which extends $\leq$ and satisfies $x\succ 1$. We fix such $\preceq$ from now on. Now consider the layer preorder $\preceq_{u}$, as defined in Remark 5.2. As we have seen, it makes $F$ into a multiplicatively Archimedean fully preordered semifield. Moreover, $\preceq_{u}$ still extends $\leq$, for the following reason. If $a\geq 1$, then the first condition in the definition of $\preceq_{u}$ in Remark 5.2 holds by the power universality of $u$ for $\leq$, and the second one holds trivially by $a\succeq 1$ and $u\succeq 1$. Hence $a\geq_{u}1$, as was to be shown. We also note $u\succ_{u}1$ as well as the important inequality
@@
By reversing all preorders and replacing $x$ by $x^{-1}$ if necessary, we can assume without loss of generality that $\preceq_{u}$ is max-tropical or max-temperate. Applying Proposition 6.7 to $(F,\preceq_{u})$ produces a $\preceq_{u}$-monotone homomorphism $\phi:F\to\mathbb{K}$ for $\mathbb{K}\in\{\mathbb{R}_{+},\mathbb{TR}_{+}\}$ with $\phi(u)>1$. The inequality (7.1) produces $\phi(x)\geq 1$. On the other hand, since $\preceq_{u}$ extends $\leq$, it is clear that $\phi$ is also $\leq$-monotone, and $\phi(u)>1$ implies that $\phi$ is nondegenerate. Therefore the assumption applies and gives $\phi(x)<1$, a contradiction.
@@
This proves the desired statement in the semifield case. It remains to reduce the general case to this one. We thus verify that the preordered semifield $F\coloneqq\operatorname{\mathsf{Frac}}(S)$ satisfies the relevant assumptions. Clearly since $\operatorname{\mathsf{Frac}}(S)/\!\sim\>=\operatorname{\mathsf{Frac}}(S/\!\sim)$, the existence of quasi-complements in $F/\!\sim$ follows from the existence of quasi-complements and quasi-inverses in $S/\!\sim$ by Lemma 3.2. Similarly, $F/\!\sim\mathop{\otimes}\mathbb{Z}$ is also a finite product of fields.
Applying the statement to $F$ then produces the desired result: the relevant inequalities $\phi\mathopen{}\mathclose{{}\left(\frac{x}{1}}\right)<\phi\mathopen{}\mathclose{{}\left(\frac{y}{1}}\right)$ and $D\mathopen{}\mathclose{{}\left(\frac{x}{1}}\right)<D\mathopen{}\mathclose{{}\left(\frac{y}{1}}\right)$ hold because every such map restricts along the homomorphism $S\to\operatorname{\mathsf{Frac}}(S)$ to a map of the corresponding type on $S$, where the assumed inequalities apply. Since a homomorphism $\phi:\operatorname{\mathsf{Frac}}(S)\to\mathbb{K}$ necessarily has trivial kernel, therefore so does its restriction to $S$. ∎
@@
    For $\mathbb{K}=\mathbb{TR}_{+}$, the optimization maps $f\mapsto\max_{s\,\in\,\operatorname{\mathsf{Newton}}(f)}\langle\beta,s\rangle$ for $\beta\in\mathbb{R}^{d}$ having components $\beta_{i}\geq 0$ not all 00.
@@
    For $\mathbb{K}=\mathbb{TR}_{+}^{\mathrm{op}}$, the optimization maps $f\mapsto\max_{s\,\in\,\operatorname{\mathsf{Newton}}(f)}\langle\beta,s\rangle$ for $\beta\in\mathbb{R}^{d}$ having components $\beta_{i}\leq 0$ not all 00.
@@ 8.  A stronger asymptotic Vergleichsstellensatz
for some $d\in\mathbb{N}$ [^13] Note that this amounts to $S/\!\sim{}\!\cong\mathbb{R}_{+}$ in the case $d=1$, and we interpret it as $S/\!\sim{}\!\cong\mathbb{B}$ for $d=0$. If this is the case, then $S/\!\sim$ obviously has quasi-complements and quasi-inverses and is such that $\operatorname{\mathsf{Frac}}(S/\!\sim)\otimes\mathbb{Z}\cong\mathbb{R}^{d}$ is a finite product of fields, so that the algebraic assumptions on $S/\!\sim$ in Theorem 7.1 are clearly satisfied.
[…]
@@
[…]
$\|\cdot\|_{1},\>\ldots,\>\|\cdot\|_{d}:S\longrightarrow\mathbb{R}_{+},$
[…]
@@
[…]
@@
The real vector space of such $\Delta$’s is exactly $\mathrm{Der}_{\mathbb{Q}}(\mathbb{R},\mathbb{R})$, which is infinite-dimensional [^18] This is by the standard fact that the module of Kähler differentials for a transcendental field extension in characteristic zero is free with basis given by a transcendence basis of the extension [1, Theorem 16.14].. Since two interchangeable derivations satisfy
@@
[…]
@@
$\mathrm{lc}_{x,y}(\phi)\coloneqq\frac{\log\frac{\phi(x)}{\phi(y)}}{\log\phi(u)},\quad\qquad\mathrm{lc}_{x,y}(D)\coloneqq\frac{D(y)-D(x)}{\|x\|_{i}},$ (8.2)
@@
Putting $d=0$ recovers Theorem 2.4 as a much simpler special case.

### v5

@@ 5. Type classification of multiplicatively Archimedean
By reversing the order and replacing $x$ by $x^{-1}$ and $r$ by 1r1-r, the second inequality reduces to the first. We therefore only prove the first.
Since the expression rx+(1r)rx+(1-r) is obviously non-strictly monotone in $r$, it is enough to prove the claim for $r=2^{-n}$ with $n\in\mathbb{N}_{>0}$, in which case it amounts to
@@
    which is equivalent to the desired inequality with $\ell^{\prime}$ in place of $\ell$. Therefore it is enough to show that if the statement holds for given $\ell\in\mathbb{N}$, then it also holds for $2\ell+1$.
    This step from $\ell$ to $2\ell+1$ works as follows,
@@
Now if $\leq$ was arctic, then we would have $x_{i}+y_{i}\sim x_{i}y_{i}+1$ by Lemma 5.15, and therefore $a+z\approx 1+z$ with $z\coloneqq\sum_{i=1}^{\ell}c_{i}(x_{i}y_{i}+1)+d$. Upon adding a quasi-complement of $z$ on both sides, we obtain further $a+n\approx 1+n$ for some $n\in\mathbb{N}$. By chaining, we can reduce to the case $n=1$ without loss of generality. But then applying the cancellation criterion of Proposition 5.13 to $1+a\leq 1+1$ shows $a\leq 1$, contradicting the initial assumption $a>1$. ∎
@@ 6. The ambient preorder
The condition $a\not\approx b$ clearly guarantees $1\,\,\cancel{\approx\approx}\,\,0$. All other required properties are also straightforward to verify, apart from the transitivity of $\leq\leq_{a,b}$. The latter is where the assumption that $F$ is a semifield (rather than a mere semiring) comes in. Indeed assuming $x\leq\leq_{a,b}y\leq\leq_{a,b}z$, we have
@@
The same definition of ambient preorder does not extend to general preordered semirings, since the transitivity may fail. For an explicit example, consider the semiring $S\coloneqq\mathbb{N}/(5\simeq 6)$, which is $\mathbb{N}$ with all numbers $\geq 5$ identified with $5$. Equip $S$ with either the trivial preorder or the total preorder inherited from $\mathbb{N}$. In either case, we have $2\leq\leq_{1,2}5\leq\leq_{1,2}1$ but $2\cancel{\leq\leq}_{1,2}1$.
@@
While the ambient preorder makes sense on any preordered semifield, we now return to the assumption that $F$ is a multiplicatively Archimedean fully preordered semifield, where the ambient preorder will facilitate the proof of our separation results. Given such an $F$, we fix an arbitrary $u\in F^{\times}$ with $u>1$. As the notation indicates, $u$ is power universal (by the definition of multiplicative Archimedeanicity). We suspect that the ambient preorder $\leq\leq_{1,u}$ is independent of the particular choice of $u$, but we have not been able to prove this so far, and we will not need it in the following. We nevertheless suppress the dependence on $u$ from our notation of the ambient preorder by writing $\leq\leq$ as shorthand for $\leq\leq_{1,u}$. In other words, we put
@@
Since reversing the original preorder $\leq$ keeps the ambient preorder $\leq\leq$ invariant, these statements hold similarly in the min-temperate and min-tropical cases, where reversing the preorder also entails that $u$ needs to be replaced by $u^{-1}$.
@@
    where the second inequality holds by power universality of $u$ and the assumed (6.1). The first inequality is strict by Lemma 5.17 in the max-temperate case, while the second inequality is strict in the max-tropical case by Lemma 5.14. Thus $x^{k}u+1<x^{k}+u$ in both cases, and we conclude $x^{k}<<1$. This implies $x<<1$ by totality of $\leq\leq$.
@@
    The definition of the ambient preorder shows that $u\geq\geq 1$ is equivalent to $u+u^{-1}\geq 2$, which we have assumed. Since $2\geq\geq 1$, this implies $v\geq\geq u\geq\geq 1$.
    For power universality, suppose first that $\leq$ is not arctic, and therefore is max-temperate or max-tropical. We then show that even $u$ itself is a power universal element for $\leq\leq$, which means that for all $x\in F^{\times}$ there is $n\in\mathbb{N}$ with $x\leq\leq u^{n}$. This latter inequality amounts to
@@
      Now for general $x$, let $y\in F$ be a quasi-complement for $x$ in $F/\!\sim$, so that $x+y\sim\ell\in\mathbb{N}_{>0}$. But what we have already shown is therefore that both $\ell$ and $\ell^{-1}(x+y)$ are upper bounded with respect to $\leq\leq$ by some power of $u$, say $u^{n}$. Hence
@@
    What we need to prove is that $4u+u^{-1}>>4$, which unfolds to
@@
By Lemma 6.6, the ambient preorder $\leq\leq$ turns $F$ into a totally preordered semifield with power universal element $v>>1$. Therefore Corollary 2.3 produces a $\leq\leq$-monotone homomorphism $\phi:F\to\mathbb{K}$ for $\mathbb{K}\in\{\mathbb{R}_{+},\mathbb{TR}_{+}\}$, where the other two cases are excluded by $v+v^{-1}>>2$.
@@
As the proof will show, the assumption on $F/\!\sim\mathop{\otimes}\mathbb{Z}$ can in fact be weakened to formal smoothness over $\mathbb{Q}$ (and probably further). We nevertheless phrase the statement in terms of the stronger assumption stating that this ring should be a finite product of fields. This is more elementary and general enough for how we will use the statement later.
@@
By quasi-complements and Lemma 6.6, the ambient preorder $\leq\leq$ has a power universal element $v>>1$ with $2<<v+v^{-1}<<2v$, so that Corollary 2.3 provides us with a $\leq\leq$-monotone homomorphism $\phi:F\to\mathbb{R}_{+}$. Consider next the multiplicative group
@@
    first for $x=u$. For $a\in\mathbb{Q}_{>0}$ it follows by (6.5). For general $a$, we can use rational approximation in the ambient preorder $\leq\leq$, which implies the claim by (c). Therefore the desired equation holds with $a=u$ for all $r$.
@@
is an ideal in $R\coloneqq(F/\!\approx)\otimes\mathbb{Z}$, namely precisely the kernel of the canonical projection homomorphism $(F/\!\approx)\otimes\mathbb{Z}\longrightarrow(F/\!\sim)\otimes\mathbb{Z}$. We have $J^{2}=0$, since $a\sim b$ and $c\sim d$ imply (ba)(dc)=0(b-a)(d-c)=0 in $R$ via Lemma 5.15 and
@@
assuming $a,c\neq 0$ without loss of generality. Hence $R$ is a square-zero extension of the quotient ring $R/J\cong\prod_{i=1}^{n}K_{i}$. Each $K_{i}$ is formally smooth over $\mathbb{Q}$ [13, Corollary 9.3.7], and therefore also their product is formally smooth over $\mathbb{Q}$. (A finite product of formally smooth algebras is formally smooth by lifting of idempotents.) In particular, the square-zero extension $R\to R/J$ is split by a ring homomorphism $\beta:R\to R/J$. In the semialgebra case, $R$ is an $\mathbb{R}$-algebra in a canonical way, and in this case we have formal smoothness over $\mathbb{R}$ for the same reason, so that we can choose $\beta$ to be $\mathbb{R}$-linear.
The universal property of $F/\!\approx\mathop{\otimes}\mathbb{Z}$ implies that $\phi:F\to\mathbb{R}_{+}$ uniquely extends to a ring homomorphism $\hat{\phi}:R\to\mathbb{R}$. In the following, we will construct a $\mathbb{Q}$-linear map
@@
    Indeed, writing a=cba=c-b for $b\sim c$ in $F^{\times}$ and plugging in the definition of $\hat{D}$ shows that this holds for all $r\in F^{\times}$. But this is nough by linearity in $r$ since $R=F^{\times}-F^{\times}$.
@@
It is interesting to ask how unique the derivation $D$ constructed in the proof is. Clearly $D$ can be replaced by any positive multiple of itself, but is there more freedom? We answer this question now.
@@ 7. A stronger catalytic Vergleichsstellensatz
We start the proof with the case where $S$ is a preordered semifield $F$ of polynomial growth. In this case, the two additional hypotheses amount to $F$ having quasi-complements and that $F/\!\sim\mathop{\otimes}\mathbb{Z}$ is a finite product of fields. Also $u\coloneqq u_{+}u_{-}^{-1}$ is a power universal element in $F$. We moreover assume $y=1$ without loss of generality. Considering this case will take the bulk of the proof; we will generalize from there in the final two paragraphs.
The goal is to prove $x\leq 1$. If $x\not\leq 1$ held, then by Theorem 2.1 we could find a total semifield preorder $\preceq$ which extends $\leq$ and satisfies $x\succ 1$. We fix such $\preceq$ from now on. Now consider the layer preorder $\preceq_{u}$, as defined in Definition 5.3. As we have seen, it turns $F$ into a multiplicatively Archimedean fully preordered semifield. Moreover, $\preceq_{u}$ still extends $\leq$, for the following reason. If $a\leq 1$ is arbitrary, then the first condition in the definition of $a\preceq_{u}1$ according to Definition 5.3 holds by the power universality of $u$ for $\leq$, and the second one holds trivially by $a\preceq 1\preceq u$. Hence $a\preceq_{u}1$, so that $\leq$ is indeed contained in $\preceq_{u}$. We also note $u\succ_{u}1$ as well as the important inequality
@@
By reversing all preorders and replacing $x$ by $x^{-1}$ if necessary, we can assume without loss of generality that $\preceq_{u}$ is max-tropical or max-temperate. Applying Proposition 6.7 to $(F,\preceq_{u})$ produces a $\preceq_{u}$-monotone homomorphism $\phi:F\to\mathbb{K}$ for $\mathbb{K}\in\{\mathbb{R}_{+},\mathbb{TR}_{+}\}$ with $\phi(u)>1$. The inequality (7.1) produces $\phi(x)\geq 1$. On the other hand, since $\preceq_{u}$ extends $\leq$, it is clear that $\phi$ is also $\leq$-monotone, and $u\geq 1$ together with $\phi(u)>1$ imply that $\phi$ is nondegenerate. Therefore the assumption applies and gives $\phi(x)<1$, a contradiction.
@@
This proves the desired statement in the semifield case. It remains to reduce the general case to this one. We thus verify that the preordered semifield $F\coloneqq\operatorname{\mathsf{Frac}}(S)$ satisfies the relevant assumptions. Clearly since $\operatorname{\mathsf{Frac}}(S)/\!\sim\>=\operatorname{\mathsf{Frac}}(S/\!\sim)$ by Lemma 3.3, the existence of quasi-complements in $F/\!\sim$ follows from the existence of quasi-complements and quasi-inverses in $S/\!\sim$ by Lemma 3.2. Similarly, $F/\!\sim\mathop{\otimes}\mathbb{Z}$ is also a finite product of fields.
Applying the statement to $F$ then produces the desired result as follows. The relevant inequalities $\phi\mathopen{}\mathclose{{\left(\frac{x}{1}}}\right)<\phi\mathopen{}\mathclose{{\left(\frac{y}{1}}}\right)$ and $D\mathopen{}\mathclose{{\left(\frac{x}{1}}}\right)<D\mathopen{}\mathclose{{\left(\frac{y}{1}}}\right)$ hold because every such map restricts along the homomorphism $S\to\operatorname{\mathsf{Frac}}(S)$ to a map of the corresponding type on $S$, where the assumed inequalities apply. Since a homomorphism $\phi:\operatorname{\mathsf{Frac}}(S)\to\mathbb{K}$ necessarily has trivial kernel, therefore so does its restriction to $S$. If $D:\operatorname{\mathsf{Frac}}(S)\to\mathbb{R}$ is a $\phi$-derivation at a degenerate homomorphism $\phi:\operatorname{\mathsf{Frac}}(S)\to\mathbb{R}_{+}$, then the restriction $D|_{S}$ is clearly also a $\phi|_{S}$-derivation, and $\phi|_{S}$ is degenerate as well since $\phi$ is degenerate. We conclude that $\frac{x}{1}\leq\frac{y}{1}$ in $\operatorname{\mathsf{Frac}}(S)$. By the definition of the preorder on $\operatorname{\mathsf{Frac}}(S)$ of (2.1), this means that there is some nonzero $a\in S$ with $ax\leq ay$, as desired. ∎
@@
    For $\mathbb{K}=\mathbb{TR}_{+}$, the optimization maps $f\mapsto\max_{s\,\in\,\operatorname{\mathsf{Newton}}(f)}\langle\beta,s\rangle$ for $\beta\in\mathbb{R}^{d}$ having components $\beta_{i}\geq 0$ not all 0.
@@
    For $\mathbb{K}=\mathbb{TR}_{+}^{\mathrm{op}}$, the optimization maps $f\mapsto\max_{s\,\in\,\operatorname{\mathsf{Newton}}(f)}\langle\beta,s\rangle$ for $\beta\in\mathbb{R}^{d}$ having components $\beta_{i}\leq 0$ not all 0.
@@ 8. A stronger asymptotic Vergleichsstellensatz
for some $d\in\mathbb{N}$. [^11] Note that this amounts to $S/\!\sim{}\!\cong\mathbb{R}_{+}$ in the case $d=1$, and we interpret it as $S/\!\sim{}\!\cong\mathbb{B}$ for $d=0$. If this is the case, then $S/\!\sim$ obviously has quasi-complements and quasi-inverses and is such that $\operatorname{\mathsf{Frac}}(S/\!\sim)\otimes\mathbb{Z}\cong\mathbb{R}^{d}$ is a finite product of fields, so that the algebraic assumptions on $S/\!\sim$ in Theorem 7.1 are clearly satisfied.
The assumption (8.1) is most conveniently formulated a little differently: we assume that $S$ is a preordered semiring equipped with a fixed surjective [^12] The surjectivity requirement can arguably be relaxed and replaced by a suitable arithmetical conditions in the image of $\|\cdot\|$ in $\mathbb{R}_{>0}^{d}\cup\{0\}$, but we have not worked this out in detail since the surjectivity holds in our applications [3, 6]. homomorphism
@@
Under these assumptions, we obtain the desired (8.1), where the isomorphism is implemented by $\|\cdot\|$ itself. The components of $\|\cdot\|$ are degenerate homomorphisms
$\|\cdot\|_{1},\>\ldots,\>\|\cdot\|_{d}:S\longrightarrow\mathbb{R}_{+}.$ (8.2)
@@
  1. (a)
    The homomorphisms (8.2) are the only degenerate homomorphisms $S\to\mathbb{R}_{+}$. [^13] This would not necessarily be the case without the surjectivity requirement on $\|\cdot\|$. For example if $d=1$, and if the image of $\|\cdot\|$ is a semiring isomorphic to $\mathbb{N}[X]$, then this fails in a particularly bad way due to the abundance of homomorphisms $\mathbb{N}[X]\to\mathbb{R}_{+}$. To see this for $d\geq 1$, use (8.1) and note that every homomorphism $\mathbb{R}_{>0}^{d}\cup\{0\}\rightarrow\mathbb{R}_{+}$ extends uniquely to a ring homomorphism $\mathbb{R}^{d}\to\mathbb{R}$, and the only such homomorphisms are the projections. For $d=0$, the statement is that there is no homomorphism $\mathbb{B}\to\mathbb{R}_{+}$ at all, which is clear.
  2. (b)
    The notation $\|\cdot\|$ is modelled after norms in functional analysis, since these sometimes have homomorphism properties when restricted to a semiring of suitably “positive” elements.
  3. (c)
    For example if $G$ is an abelian group and $\mathbb{R}_{+}[G]$ the associated group semialgebra, then the $\ell^{1}$-norm is a semiring homomorphism $\mathbb{R}_{+}[G]\to\mathbb{R}_{+}$, making it an instance of the above with $d=1$.
@@
The real vector space of such $\Delta$’s is exactly $\mathrm{Der}_{\mathbb{Q}}(\mathbb{R},\mathbb{R})$, which is infinite-dimensional [^15] This is by the standard fact that the module of Kähler differentials for a transcendental field extension in characteristic zero is free with basis given by a transcendence basis of the extension [2, Theorem 16.14].. Since two interchangeable derivations satisfy
@@
$\sqcup\bigsqcup_{i=1}^{d}\{\text{{monotone }}\|\cdot\|_{i}\text{{-derivations }}S\to\mathbb{R}\text{{ with }}D(u)=1\text{{ modulo interchangeability}}\}.$
@@
$\mathrm{lc}_{x,y}(\phi)\coloneqq\frac{\log\frac{\phi(y)}{\phi(x)}}{\log\phi(u)},\quad\qquad\mathrm{lc}_{x,y}(D)\coloneqq\frac{D(y)-D(x)}{\|x\|_{i}},$ (8.3)
@@
      For every $i=1,\ldots,d$ and monotone $\|\cdot\|_{i}$-derivation [^16] Recall from (8.2) that we write $\|\cdot\|_{i}$ for the $i$-th component of $\|\cdot\|$. $D:S\to\mathbb{R}$ with $D(u)=1$,
@@
Putting $d=0$ lets us recover Theorem 2.4 as a much simpler special case, since $\mathbb{R}_{>0}^{d}\cup\{0\}=\mathbb{B}$ is the Boolean semiring, the derivations therefore do not appear at all, and by the assumed $1>0$ only spectral points with $\mathbb{K}\in\{\mathbb{R}_{+},\mathbb{TR}_{+}\}$ can occur.
@@
The implication from (b) to (a) is again simple. Indeed if (b) holds, then we treat the two kinds of spectral points separately:
  • $\triangleright$
    For $\phi:S\to\mathbb{K}$ with $\mathbb{K}\in\{\mathbb{R}_{+},\mathbb{R}_{+}^{\mathrm{op}},\mathbb{TR}_{+},\mathbb{TR}_{+}^{\mathrm{op}}\}$ a nondegenerate monotone homomorphism with trivial kernel, we have $\phi(u)>1$ by power universality of $u$, as well as
    $\phi(x)^{n}\leq\phi(u)^{\lfloor\varepsilon n\rfloor}\phi(y)^{n}$
    by monotonicity and multiplicativity of $\phi$. By taking $n$-th roots and letting $\varepsilon\to 0$, we obtain the desired $\phi(x)\leq\phi(y)$.
  • $\triangleright$
    For $i=1,\ldots,d$ and $D:S\to\mathbb{R}$ a monotone $\|\cdot\|_{i}$-derivation with $D(u)=1$, we have:
    $nD(x)=D(x^{n})\leq D(u^{\lfloor\varepsilon n\rfloor}y^{n})=\lfloor\varepsilon n\rfloor D(u)+D(y^{n})=\lfloor\varepsilon n\rfloor+nD(y).$
    Dividing by $n$ and letting $\varepsilon\to 0$ gives the desired $D(x)\leq D(y)$.
For the other implications, note first that the assumed $S/\!\sim{}\!\cong\mathbb{R}_{>0}^{d}\cup\{0\}$ implies that $S$ is a zerosumfree semidomain, so that Theorem 7.1 applies. Thus if (a) holds with strict inequalities, then there is nonzero $b\in S$ with $bx\leq by$. Choosing any $c\in S$ with $\|c\|=\|b\|^{-1}$ by the surjectivity of $\|\cdot\|$ and taking $a\coloneqq bc$ results in $\|a\|=1$, and therefore we can achieve
$ax\leq ay,\qquad a\sim 1.$